What would the pressure be if the volume of the syringe was increased to 40.0 mL?

And what is Boyle's law equation? Is it it K=PV or K=P/V
Answer #1

Boyle's law equation is K = PV.

Since, P1*V1 = P2*V2 (According to Boyle's law) If the initial pressure is p1 and initial volume is v1 (in mL), we can write

p1*v1 = p2*(40 mL)

so the p2 will be = (p1*v1)/40